OpenStudy (anonymous):

Martha paid $42.40 for her sister’s birthday present. This included 6% sales tax. What was the cost of the gift before tax? Would this be $40.00

OpenStudy (anonymous):

Write an equation and solve. Round to the nearest hundredth where necessary. 19 is what percent of 40?

OpenStudy (anonymous):

21 is 35% of what number?

OpenStudy (anonymous):

frist one is correct

OpenStudy (anonymous):

Thankyou!

OpenStudy (anonymous):

60 for the 2nd one because 35% of 60 is 21 :)

OpenStudy (anonymous):

What about the others?

OpenStudy (anonymous):

Amount Martha paid = $ 42.40 Sales tax = 6% So, the sales tax is = 6/100*42.40 = 2.544 Amount without sales tax = $ 42.40-2.544 = $ 39.856 = $ 40. Second Question 35/100*x = 21 35x/100 = 21 Cross multiply 35x = 21*100 x = 21*100/35 = 60

OpenStudy (anonymous):

last one is 47.5%

OpenStudy (anonymous):

Sorry I meant the last question 35/100*x = 21 35x/100 = 21 Cross multiply 35x = 21*100 x = 21*100/35 = 60

OpenStudy (anonymous):

For the second question, we take the % as x% so, x/100*40 = 19 Cross multiply. 40x/100=19 40x=19*100 x = 1900/40 = 47.5%

OpenStudy (anonymous):

Thankyou!!!

OpenStudy (anonymous):

=)